# Learning topological states from randomized measurements using variational tensor network tomography

### v3

@@ F.1 Perturbed surface code
where $\mathcal{S}=\{S_{Z,\,l},S_{X,\,l},\,l=1,\cdots,\frac{n-1}{2}\}$ is the set of $n-1$ stabilizers around a square, as illustrated in Fig. 2(a). To construct the Hamiltonian for an $(L_{x},L_{y})=(3,3)$ system, the following lines create an object called PhysicalSystem and generate its matrix product operator representation.
@@ F.2 Ruby-lattice quantum spin liquid
[…]
@@ G.1 Optimization scheme via gradient descent
We approach the optimization of matrix product state (MPS) tensors by updating all tensors simultaneously. While the target states we are learning are real states, we use MPS tensors with complex variables because we have found that optimization over complex tensors is more robust, as also observed in Ref.  [27]. The minimization of our loss function involves the logarithm of sums over expectation values of $U_{i}^{\dagger}\rvert b_{i}\rangle\!\langle b_{i}\lvert U_{i}$ defined by the measured unitary-bit-strings $D=\{(U_{i},b_{i})\}_{i=1}^{N}$,
@@ G.2 Model selection
Supplementary Figure 3: Model selection by comparing training and test losses for a surface code with different system sizes and measurement schemes. On the dashed lines, the training NLL $=$ test NLL; ideally, the model should be selected along the dashed line and with a small training NLL. The models which have small NLLs and are close to the dashed line yield high fidelities (darker colors).

### v4

@@ F.1 Perturbed surface code
where $\mathcal{S}=\{S_{Z,\,l},S_{X,\,l},\,l=1,\cdots,\frac{n-1}{2}\}$ is the set of n1n-1 stabilizers around a square, as illustrated in Fig. 2(a). To construct the Hamiltonian for an $(L_{x},L_{y})=(3,3)$ system, the following lines create an object called PhysicalSystem and generate its matrix product operator representation.
@@ F.2 Ruby-lattice quantum spin liquid
Denoting by $\mbox{$|b\rangle$}=\mbox{$|\pm\rangle$}$ the eigenstate of the Pauli $Z$ operator in the computational basis (i.e., $Z\mbox{$|b\rangle$}=b\mbox{$|b\rangle$}$), we work with a sign convention such that $b=+1$ (1-1) corresponds to an atom being in the Rydberg (ground) state.
@@ G.1 Optimization scheme via gradient descent
We approach the optimization of matrix product state (MPS) tensors by updating all tensors simultaneously. While the target states we are learning are real states, we use MPS tensors with complex variables because we have found that optimization over complex tensors is more robust, as also observed in Ref.  [27]. The minimization of our loss function involves the logarithm of sums over expectation values of $U_{i}^{\dagger}\rvert b_{i}\rangle\!\langle b_{i}\lvert U_{i}$ defined by the measured unitary-bitstrings $D=\{(U_{i},b_{i})\}_{i=1}^{N}$,
@@ G.2 Model selection
Supplementary Figure 3: Model selection by comparing training and test losses for a surface code with different system sizes and measurement schemes. On the dashed lines, the training NLL == test NLL; ideally, the model should be selected along the dashed line and with a small training NLL. The models which have small NLLs and are close to the dashed line yield high fidelities (darker colors).
APPENDIX H Measurement noise
In this section, we discuss the effect of measurement noise in our tomographic protocol. As we have emphasized in the main text, the motivations for choosing our random-$XZ$ measurements are twofold: 1) such measurements are sufficient and tomographically complete for the real pure states of interests; 2) these measurements only require single-qubit rotations before measurements, suggesting their experimental feasibility. Nevertheless, there are imperfections in realistic measurements even in the most well-developed systems. A rigorous treatment of such errors is left for future work. Here, we numerically study how the fidelity of the reconstructed state is affected by measurement noise.
Rydberg atom arrays have seen significant recent progress, achieving impressive gate fidelities. In digital simulations, qubits are typically encoded in the hyperfine states $\{\mbox{$|0\rangle$},\mbox{$|1\rangle$}\}$ of 87Rb atoms. In this setting, the most significant errors originate from many-qubit gates such as the controlled-phase entangling gate. Currently, the state-of-the-art two-qubit entangling gate fidelity is $99.5\%$ and single-qubit gate fidelity reaches $99.9\%$ [64].
In contrast, the states studied in this work are prepared through analog evolution, where interactions are mediated by the strong repulsive forces between atoms in the Rydberg state $|r\rangle$ in a two-level system of $\{\mbox{$|1\rangle$},\mbox{$|r\rangle$}\}$. Since the Rydberg state is short-lived and hard to measure directly, a “coherent mapping” between the Rydberg and hyperfine manifolds is necessary. The mapping can be achieved by first mapping the hyperfine excited state to the ground state $\mbox{$|1\rangle$}\rightarrow\mbox{$|0\rangle$}$ via a Raman $\pi$ pulse, and then mapping the Rydberg state to the hyperfine excited state $\mbox{$|r\rangle$}\rightarrow\mbox{$|1\rangle$}$ via a Rydberg $\pi$ pulse. Beyond the imperfections of the pulses (which could be neglected), there are additional sources of errors from the coherent mapping. Such a protocol is analogous to a hybrid simulation presented in Fig. 4 of Ref. [12], wherein a state is prepared via analog simulation and thereafter manipulated by digital evolution. The main sources of errors during such a coherent mapping are [12] the following:
  1. 1.
    Failure of the Rydberg $\pi$ pulse. With a small probability, two neighboring 87Rb atoms both occupy the Rydberg state $|r\rangle$, and such a violation of the Rydberg blockade would dramatically shift the energy level. When this occurs, the pair of Rydberg states cannot be converted to the hyperfine states, leading to an incomplete sample.
  2. 2.
    Off-resonance of Rydberg $\pi$ pulse. Even without such a violation, the long-range interaction between the Rydberg states would slightly detune the Rydberg $\pi$ pulse, leading to a different phase factor in the mapped state.
  3. 3.
    Dephasing during Raman $\pi$ pulse (Doppler shifts). When the atoms are idling, their atomic motions lead to off-resonance of the Raman pulse, called the Doppler shift. As the Raman pulse has a small frequency, and is therefore slow compared to the Rydberg pulse, the atoms can accumulate a more substantial random phase.
In summary, because a single-qubit gate of an analog-prepared state requires coherent mappings between the Rydberg state and the hyperfine state, additional errors during such mappings contribute to the errors in these single-qubit gates. It is therefore important to carefully consider these measurement errors in future work. As a first step, we consider a simple bit-flip noise model where $|0\rangle$ and $|1\rangle$ are misclassified with each other with a probability of $\xi$. We compare the fidelity achieved from training with a noisy dataset with probability $\xi$ for the $3\times 3$ surface code state in Supplementary Fig. 4. As expected, higher noise levels lead to a reduced fidelity. The infidelity scales continuously with NN—even in the presence of noise—at an initial stage of training, and exhibits a plateau with a different scaling rate after a certain threshold number of measurements. Such a scaling is worth exploring systematically in future work.
Supplementary Figure 4: Robustness of the tomographic protocol to measurement noise. We study the effect of measurement noise modeled by a symmetric bit-flip channel, where measurement outcomes $|0\rangle$ and $|1\rangle$ are flipped with probability $\xi$. The plot shows the infidelity 1F1-F of the reconstructed $3\times 3$ surface code state as a function of the number of samples $N$, for varying noise rate $\xi$. Results are shown for the random-$XZ$ measurements.